Find the order and degree of the following differential equations

a)      D^y/dx^3+6d^2y/dx+11dy/dx+6y=0

b)     (D^3y/dx^3)^2-3d^2y/dx^2+4y=0

c)      (1-x^2)dy/dx-xy=1

d)     2d^2y/dx^2-3dy/dx+y=0


1
Expert's answer
2022-04-01T11:55:38-0400

The ORDER of an ODE is the highest derivative in the equation


The DEGREE of an ODE is the power of the highest derivative in the equation



(a) Highest derivative is "\frac{d³y}{dx³}". Thus, the order is 3.

The power of the highest derivative is 1. Hence, the degree is 1





(b) Highest derivative is "\frac{d³y}{dx³}" . Thus, the order is 3.

The power of the highest derivative is 2. Hence, the degree is 2





(c) Highest derivative is "\frac{dy}{dx}" . Thus, the order is 1.

The power of the highest derivative is 1. Hence, the degree is 1.




(d) Highest derivative is "\frac{d²y}{dx²}" . Thus, the order is 2.

The power of the highest derivative is 1. Hence, the degree is 1